a) based on the self-ionized equilibrium for water, calculate the hydroxide ion concentration [OH-] when the hydronium ion concentration is [H3O+]=1.67x10^-4

b)calculate the pH of the solution

c)is the solution described above acidic, basic, or neutral? Explain
1
Expert's answer
2015-11-16T09:50:55-0500
Answer:
a) Kw=[H+][OH-]=10-14
[OH-]=10-14/1.67∙10-4=5.99∙10-11
b) pH=-lg[H+]
pH=-lg[1.67∙10-4]=3.78
c) The solution is acidic because pH<7